# Vision-Centric 4D Occupancy Forecasting and Planning via Implicit Residual World Models

### v2

@@ Abstract
End-to-end autonomous driving systems increasingly rely on vision-centric world models to understand and predict their environment. However, a common ineffectiveness in these models is the full reconstruction of future scenes, which expends significant capacity on redundantly modeling static backgrounds. To address this, we propose IR-WM, an Implicit Residual World Model that focuses on modeling the current state and evolution of the world. IR-WM first establishes a robust bird’s-eye-view representation of the current state from the visual observation. It then leverages the BEV features from the previous timestep as a strong temporal prior and predicts only the “residual”, i.e., the changes conditioned on the ego-vehicle’s actions and scene context. To alleviate error accumulation over time, we further apply an alignment module to calibrate semantic and dynamic misalignments. Moreover, we investigate different forecasting–planning coupling schemes and demonstrate that the implicit future state generated by world models substantially improves planning accuracy. On the nuScenes benchmark, IR-WM achieves top performance in both 4D occupancy forecasting and trajectory planning.
@@ I Introduction
End-to-end autonomous driving models, which directly map raw sensor data to planning outputs, have become a dominant paradigm [1, 2, 3, 4]. At its core, safe decision-making in autonomous driving hinges on the precise understanding of current and past status, along with the reliable anticipation of future events. Consequently, action-driven world models have emerged as a promising approach that bridges agent intentions and environmental dynamics through action-guided world generation and forecast-driven planning, thereby enabling more reliable decision-making.
@@ II-A Driving World Models
World models aim to simulate future environment states given past observations and actions. Image-based approaches [11, 12, 13, 14, 15, 16, 17, 18] generate future driving videos using diffusion or transformer architectures, often conditioned on actions, BEV layouts, or HD maps to enhance realism. In contrast, volume-based methods [19, 20, 21, 7, 22, 23] directly model 3D structures such as occupancy or point clouds, enabling spatially consistent future prediction. Unlike recent approaches, we construct a residual BEV world model with feature alignment, guided by 4D occupancy forecasting, which balances model capacity to effectively capture both the current state and its future changes. This design alleviates error accumulation, leverages vision-centric state modeling as a foundation for accurate future prediction, and strengthens downstream planning.
@@ II-B End-to-end Driving
End-to-end models have emerged as a paradigm shift in autonomous driving, directly mapping sensor inputs to future trajectories [3, 4]. Early approaches such as P3 [24] and ST-P3 [25] incorporated differentiable occupancy to ensure safe planning. More recent frameworks leverage BEV perception as an intermediate representation, including UniAD [3], VAD [4], and GraphAD [26]. Further advances focus on disentangled scene representations [27], reducing annotation costs [28, 5], or improving efficiency through parallel or sparse architectures [29, 30, 31]. Our work explores the impact of implicit representation generated by world models on end-to-end planning.
@@ III-B2 Future Forecasting with Streaming Memory
Feature Alignment. To obtain the final BEV features $\mathbf{S}^{bev}_{t}$, we add the predicted feature differences $\Delta\mathbf{S}^{bev}_{t}$ with previous BEV features $\mathbf{S}^{bev}_{t-1}$. Then we apply a feature alignment module further to enhance both the semantic richness and dynamic understanding, providing aligned features for the next rollout to alleviate potential error accumulation. Inspired by AdaNorm [46] and [8], instead of just adding the information, we modulate the features by dynamically generating the scale and shift parameters ($\gamma,\beta$) of layer normalization. The scale and shift parameters are generated from the predicted semantic occupancy maps $\mathbf{O}^{align}_{t}$ and the ego-motion transformations $\delta\tau_{t}$. The procedure can be formulated as:
@@
$\mathbf{\hat{S}}^{bev}_{t}=\gamma(\mathbf{O}^{align}_{t})\cdot\mathrm{LN}(\mathbf{S}^{bev}_{t})+\beta(\mathbf{O}^{align}_{t})+$ (7)
$\gamma(\delta\tau_{t})\cdot\mathrm{LN}(\mathbf{S}^{bev}_{t})+\beta(\delta\tau_{t})$ (8)
@@
Figure 3: The detailed architecture of the planning head. “Adapter” is a stack of convolutions.
@@
Planning Head. As depicted in Fig. 3, for planning, a learnable plan query predicts the ego translation $\delta\tau_{t+1}$, updating the pose as $\tau_{t+1}=\tau_{t}+\delta\tau_{t+1}$. At timestamp $t$, the plan query is concatenated with the command embedding and passed to a transformer decoder that performs cross-attention over $\textbf{S}_{t}^{bev}+\rm{Adapter}(\textbf{S}_{t}^{bev})$ to capture fine-grained scene context and attend to potential dynamic obstacles, followed by a feed-forward network; an MLP then outputs $\delta\tau_{t+1}$. The adapter is a stack of convolutions for feature enhancement.
@@ III-C Training Objective
Figure 5: Qualitative results for 4D occupancy forecasting and planning. IR-WM produces reliable predictions for both static and dynamic objects across different times of day.
@@ IV-A3 Implementation Details
Method $\text{mIoU}_{c}$
$\text{mIoU}_{f}$
@@ IV-B3 End-to-End Planning
Variants $\text{mIoU}_{c}$
$\text{mIoU}_{f}$
(2.5 s)
$\tilde{\text{mIoU}}_{f}$ L2 (m) $\downarrow$
Latency
@@
TABLE VI: Ablations on auxiliary 3D flow supervision. We provide Inflated GMO Forecasting results on nuScenes validation.
Method $\text{IoU}_{c}$ $\text{IoU}_{f}$ (2 s) $\tilde{\text{mIoU}}_{f}$
IR-WM 40.80 37.20 38.20
w/ Flow 41.20 37.80 38.80
@@ IV-D Analysis
TABLE VII: Ablations on the temporal self-supervision (TSS) on BEV features. We provide Fine-Grained GMO and GSO Forecasting results on nuScenes-Occupancy validation.
@@
TABLE VIII: Ablations on the action conditioning. We provide Fine-Grained MMO and MSO Forecasting results on nuScenes-Occupancy validation.
@@
TABLE IX: Ablations on the feature alignment (FA). We provide per-class ${\text{IoU}}_{f}$ results on nuScenes-Occupancy validation.
@@
TABLE X: Ablations on the BEV feature resolution. We provide Fine-grained GMO Forecasting results on nuScenes-Occupancy validation.
@@
Impact of Condition Interface and Feature Alignment. We evaluate the effect of action conditioning and the alignment module. The detailed results are summarized in Table VIII. Using addition for conditioning achieves better performance compared to cross-attention while reducing the parameters. Moreover, as shown in Table IX, removing the alignment module leads to a drop in $\text{IoU}_{f}$ for typical static objects, indicating that proper alignment of predicted BEV features can slightly alleviate error accumulation over time.
Exploration on Temporal Self-Supervision. We further evaluate temporal self-supervision (TSS) on BEV features. During training, the scene encoder $\mathcal{E}$ extracts BEV features $\mathbf{{S}}^{obs}_{1:f-1}$ from future multiview images $\mathbf{I}_{1:f-1}$, and an L2 loss is applied to supervise the predicted BEV features $\mathbf{S}^{bev}_{1:f-1}$. As shown in Table VII, TSS provides only a marginal regularization effect on occupancy prediction. This indicates that our implicit residual world model, guided by 4D occupancy, already enforces strong temporal consistency, leaving limited benefit for additional temporal constraints.
Impact of 3D flow supervision. We assess the benefit of auxiliary 3D flow supervision. Results in Table VI show that incorporating flow prediction improves both current and future IoU, confirming that motion-aware supervision enhances the model’s ability to reason about future occupancy. However, obtaining dense 3D flow labels is costly, so our method does not rely on them, even though the ablation verifies their effectiveness.
Impact of Different BEV Feature Resolutions. Finally, we compare different BEV feature resolutions. Results in Table X showcase that our method consistently outperforms the Drive-OccWorld baseline under the same resolution, demonstrating the effectiveness of our implicit residual world model. Moreover, higher spatial resolution ($\times200$) provides a significant advantage over coarser features ($\times128$), underlining the importance of fine-grained BEV representation for accurate occupancy prediction.

### v3

@@ Abstract
End-to-end autonomous driving systems increasingly rely on vision-centric world models to understand and predict their environment. However, a common ineffectiveness in these models is the full reconstruction of future scenes, which expends significant capacity on redundantly modeling static backgrounds. To address this, we propose IR-WM, an Implicit Residual World Model that focuses on modeling the current state and evolution of the world. IR-WM first establishes a robust bird’s-eye-view representation of the current state from the visual observation. It then leverages the BEV features from the previous timestep as a strong temporal prior and predicts only the “residual”, i.e., the changes conditioned on the ego-vehicle’s actions and scene context. To alleviate error accumulation over time, we further apply an alignment module to calibrate semantic and dynamic misalignments. Moreover, we investigate different forecasting–planning coupling schemes and demonstrate that the implicit future state generated by world models substantially improves planning accuracy. On the nuScenes benchmark, IR-WM achieves top performance in both 4D occupancy forecasting and trajectory planning. Codes are available at https://github.com/yuyang-cloud/Drive-OccWorld
@@ I Introduction
End-to-end autonomous driving models, which directly map raw sensor data to planning outputs, have become a dominant paradigm [9, 11]. At its core, safe decision-making in autonomous driving hinges on the precise understanding of current and past status, along with the reliable anticipation of future events. Consequently, action-driven world models have emerged as a promising approach that bridges agent intentions and environmental dynamics through action-guided world generation and forecast-driven planning, thereby enabling more reliable decision-making.
@@ II-A Driving World Models
World models aim to simulate future environment states given past observations and actions. Image-based approaches [6, 32, 16, 42, 13, 22, 43] generate future driving videos using diffusion or transformer architectures, often conditioned on actions, BEV layouts, or HD maps to enhance realism. In contrast, volume-based methods [51, 54, 31, 25, 37, 45] directly model 3D structures such as occupancy or point clouds, enabling spatially consistent future prediction. Unlike recent approaches, we construct a residual BEV world model with feature alignment, guided by 4D occupancy forecasting, which balances model capacity to effectively capture both the current state and its future changes. This design alleviates error accumulation, leverages vision-centric state modeling as a foundation for accurate future prediction, and strengthens downstream planning.
@@ II-B End-to-end Driving
End-to-end models have emerged as a paradigm shift in autonomous driving, directly mapping sensor inputs to future trajectories [9, 11]. Early approaches such as P3 [27] and ST-P3 [8] incorporated differentiable occupancy to ensure safe planning. More recent frameworks leverage BEV perception as an intermediate representation, including UniAD [9], VAD [11], and GraphAD [52]. Further advances focus on disentangled scene representations [4], reducing annotation costs [5, 14], or improving efficiency through parallel or sparse architectures [36, 50]. Our work explores the impact of implicit representation generated by world models on end-to-end planning.
@@ III-B2 Future Forecasting with Streaming Memory
Feature Alignment. To obtain the final BEV features $\mathbf{S}^{bev}_{t}$, we add the predicted feature differences $\Delta\mathbf{S}^{bev}_{t}$ with previous BEV features $\mathbf{S}^{bev}_{t-1}$. Then we apply a feature alignment module further to enhance both the semantic richness and dynamic understanding, providing aligned features for the next rollout to alleviate potential error accumulation. Inspired by AdaNorm [38] and [46], instead of just adding the information, we modulate the features by dynamically generating the scale and shift parameters ($\gamma,\beta$) of layer normalization. The scale and shift parameters $(\gamma_{s},\beta_{s})$ for semantics and $(\gamma_{e},\beta_{e})$ for dynamics are generated by separate MLPs from the predicted semantic occupancy maps $\mathbf{O}^{align}_{t}$ and ego-motion transformations $\delta\tau_{t}$, respectively. The procedure can be formulated as:
@@
$\mathbf{\hat{S}}^{bev}_{t}=\gamma_{s}(\mathbf{O}^{align}_{t})\cdot\mathrm{LN}(\mathbf{S}^{bev}_{t})+\beta_{s}(\mathbf{O}^{align}_{t})+$ (7)
$\gamma_{e}(\delta\tau_{t})\cdot\mathrm{LN}(\mathbf{S}^{bev}_{t})+\beta_{e}(\delta\tau_{t})$ (8)
@@
Figure 3: The detailed architecture of the planning head.
@@
Planning Head. As shown in Fig. 3, for planning, a learnable plan query predicts the ego translation $\delta\tau_{t+1}$, updating the pose as $\tau_{t+1}=\tau_{t}+\delta\tau_{t+1}$. At timestamp $t$, the plan query is concatenated with the command embedding and passed to a transformer decoder that performs cross-attention over $\textbf{S}_{t}^{bev}+\rm{Adapter}(\textbf{S}_{t}^{bev})$ to capture fine-grained scene context and attend to potential dynamic obstacles, followed by a feed-forward network; an MLP then outputs $\delta\tau_{t+1}$. The adapter is a stack of convolutions for feature enhancement.
@@ III-C Training Objective
Figure 5: Qualitative results for 4D occupancy forecasting and planning show that our method yields reliable predictions for static and dynamic objects at different times of day.
@@ IV-A3 Implementation Details
Method $\text{mIoU}_{c}$ $\text{mIoU}_{f}$ (2.5s)
@@ IV-B3 End-to-End Planning
Variants $\text{mIoU}_{c}$ $\text{mIoU}_{f}$ (2.5 s) $\tilde{\text{mIoU}}_{f}$ L2 (m) $\downarrow$ Latency (ms)
@@ IV-D Analysis
TABLE VI: Ablations on the temporal self-supervision (TSS) on BEV features. We provide Fine-Grained GMO and GSO Forecasting results on nuScenes-Occupancy validation.
@@
TABLE VII: Ablations on the action conditioning. We provide Fine-Grained MMO and MSO Forecasting results on nuScenes-Occupancy validation.
@@
TABLE VIII: Ablations on the feature alignment (FA). We provide per-class ${\text{IoU}}_{f}$ results on nuScenes-Occupancy validation.
@@
TABLE IX: Ablations on the BEV feature resolution. We provide Fine-grained GMO Forecasting results on nuScenes-Occupancy validation.
@@
Impact of Condition Interface and Feature Alignment. We evaluate the effect of action conditioning and the alignment module. The detailed results are summarized in Table VII. Using addition for conditioning achieves better performance compared to cross-attention while reducing the parameters. Moreover, as shown in Table VIII, removing the alignment module leads to a drop in $\text{IoU}_{f}$ for typical static objects, indicating that proper alignment of predicted BEV features can slightly alleviate error accumulation over time.
Exploration on Temporal Self-Supervision. We further evaluate temporal self-supervision (TSS) on BEV features. During training, the scene encoder $\mathcal{E}$ extracts BEV features $\mathbf{{S}}^{obs}_{1:f-1}$ from future multiview images $\mathbf{I}_{1:f-1}$, and an L2 loss is applied to supervise the predicted BEV features $\mathbf{S}^{bev}_{1:f-1}$. As shown in Table VI, TSS provides only a marginal regularization effect on occupancy prediction. This indicates that our implicit residual world model, guided by 4D occupancy, already enforces strong temporal consistency, leaving limited benefit for additional temporal constraints.
Impact of Different BEV Feature Resolutions. Finally, we compare different BEV feature resolutions. Results in Table IX showcase that our method consistently outperforms the Drive-OccWorld baseline under the same resolution, demonstrating the effectiveness of our implicit residual world model. Moreover, higher spatial resolution ($\times200$) provides a significant advantage over coarser features ($\times128$), underlining the importance of fine-grained BEV representation for accurate occupancy prediction.